OpenStudy (tommo_lcfc):

I am not asking these questions because I need some help, I am asking these questions because I have been helping all of you with your problems and now you need to return the favour. I am going to set all of you 6 questions on general mathematics. Let's see how many you can get correct! 1) Solve x^2+5x+6 = 0 by factorising. 2) Solve x^2 - 2x + 2 = 0 by completing the square. 3) Solve x^2 - 6x + 13 = 0 by using the quadratic formula. 4) If y = 3x^2, find dy/dx. Also integrate this. 5) If y = 3x + 4 and x = 10, find y. 6) Expand (x-3)(x+3). What is this expression an example of?

OpenStudy (tommo_lcfc):

I will be giving medals out to everyone who gets the answers correct.

OpenStudy (callisto):

1. x^2+5x+6 = 0 (x+3)(x+2) =0 x=-2 or x=-3

OpenStudy (callisto):

2. x^2 - 2x + 2 = 0 (x^2 -2x+1) +1 =0 (x-1)^2 =-1 (x-1) = i or -i x = i+1 or x = 1-i

OpenStudy (callisto):

3. x^2 - 6x + 13 = 0 \[x=[-(-6) \pm \sqrt{(-6)^{2}-4(1)(13)}] / 2\]\[x=(6 \pm \sqrt{(-16} )/ 2\]\[x=(3 \pm 2i)\]

OpenStudy (callisto):

4. dy/dx = d/dx (3x^2) = 6x \[\int\limits 3x^2 dx = 3 (1/3) x^3 +C = x^3 +C\]

OpenStudy (callisto):

5. y = 3x + 4 and x = 10 Put x=10 into y = 3x+4 y= 3(10) +4 = 34

OpenStudy (callisto):

6. (x-3)(x+3) = x^2 - 9 It is difference of 2 square

OpenStudy (callisto):

Please mark the answers !

OpenStudy (tommo_lcfc):

I've already given you the medals, apparently! I knew I had given one medal but didn't know I had given the medal for every answer you wrote down! You have got everything right. I need to give you a harder quiz!

OpenStudy (callisto):

thanks